# It’s Giving Major Satisfaction: Why Fairness Matters for Developers

### v1

@@
Abstract
Software practitioners often face unfairness in their work, such as unequal recognition of contributions, gender bias, and unclear criteria for performance reviews. While the link between fairness and job satisfaction has been established in other fields, its relevance to software professionals remains underexplored. This study aims to examine how fairness perceptions relate to job satisfaction among software practitioners, focusing on both general trends and demographic-specific differences. We conducted an online survey of 108 software practitioners, followed by ordinal logistic regression to analyze the relationship between fairness perceptions and job satisfaction in software engineering contexts, with moderation analysis examining how this relationship varies across demographic groups.
@@ Abstract
Our findings indicate that all four fairness dimensions, distributive, procedural, interpersonal, and informational, significantly affect both overall job satisfaction and satisfaction with job security. Among these, interpersonal fairness has the biggest impact, being more than twice as influential on overall job satisfaction. The relationship between fairness perceptions and job satisfaction is notably stronger for female, ethnically underrepresented, less experienced practitioners, and those with work limitations. Fairness in authorship emerged as an important factor for job satisfaction collectively, while fairness in policy implementation, high-demand situations, and working hours particularly impacted specific demographic groups. This study highlights the unique role of fairness in software engineering, offering strategies for organizations to promote fair practices and targeted approaches specific for certain demographic groups.
Index Terms:
Fairness, human aspects, social aspects, software engineering
I Introduction
@@
In the field of Software Engineering (SE), practitioners frequently encounter fairness problems in various aspects of their work, such as problems in recognition of contributions [1], gender bias [2], and ethnicity bias [3]. A study by Sesari et al.[4] explored these fairness problems by identifying different contexts in which they arise, using Colquitt’s well-established organizational fairness framework [5]. By leveraging this framework, the authors were able to pinpoint specific areas within SE where fairness issues are most prevalent. The top five areas identified were income, treatment, allocation of work, recruitment, and working hours [4].
@@
Leveraging the same framework, German et al. developed a framework explaining fairness in the modern code review process to identify strengths and potential areas for improvement [6]. For example, distributive fairness, which includes principles like equity [5], suggests that decisions about how work is evaluated and rewarded should consider both the person who made the contribution and the quality of that contribution. German et al. further studied fairness problems among developers involved in the OpenStack project by surveying them about the code review process [6]. The survey results showed that many developers were concerned about fairness, especially regarding issues like equity, equal treatment, consistency, and preventing bias [6]. However, while these problems have been identified, the impact of facing such fairness issues on software developers remains unclear.
To better understand the broader implications of fairness in the workplace, Colquitt conducted a meta-analytic review of organizational justice literature, which included 120 separate meta-analyses of 183 empirical studies [7]. This comprehensive review examined the effects of organizational fairness—encompassing distributive, procedural, interpersonal, and informational fairness—on several outcomes, including job satisfaction. Colquitt’s findings demonstrate a significant link between perceptions of fairness and job satisfaction across diverse contexts.
@@
For example, in the healthcare sector, fairness in managerial practices has been shown to significantly influence nurse retention and job satisfaction [8]. In the manufacturing industry, procedural justice has been found to enhance employee commitment and reduce turnover intentions [9]. Similarly, in educational settings, distributive fairness related to workload and rewards is a key predictor of teacher satisfaction [10]. Given these findings across various fields, an important question arises: does the observed link between fairness and job satisfaction apply in the context of SE?
@@
In SE literature, enhanced job satisfaction among developers is known to improve employee attraction, retention, and well-being [11]. Furthermore, many researchers and companies assume that more satisfied developers tend to be more productive [12]. Storey et al. developed a theory that links higher job satisfaction among software developers to increased perceived productivity [12]. In their study, they also identified various social and technical factors that significantly impact job satisfaction, such as appreciation and rewards, work culture, and work life balance.
@@
Building on this work, we are interested in exploring perception of fairness as another potential factor that could shape job satisfaction. Similar to Storey et al.’s approach in identifying key factors that influence job satisfaction, our study seeks to investigate whether and how perceptions of fairness relate to job satisfaction among software practitioners, both as a collective, and from different backgrounds.
II Research Questions
@@
To address whether and how perceptions of fairness relate to job satisfaction, we conducted an online survey, receiving a total of 108 responses. We used the 20-item fairness scale developed by Colquitt [5] to measure fairness perceptions across four dimensions: distributive, procedural, interpersonal, and informational fairness. Job satisfaction was measured using 15 satisfaction factors inspired by Storey et al. [12].
@@
This observation aligns with existing literature that highlights demographic disparities in the SE industry, which is predominantly young, White, and male. Studies have shown that minority groups often face unique challenges. For example, Trinkenreich et al. identified that women in the software industry encounter issues like work-life balance, the glass ceiling, and sexism [13]. Canedo et al. found that women are underrepresented in software development and open source projects, often due to male-dominated workplace conditions [2]. Furthermore, Nadri et al. reported that less than 10% of contributions to open source projects come from visibly Non-White developers [3]. Similarly, Baltes et al. discuss age-related biases, noting that older developers are often perceived as less adaptable, which can limit their career progression [14].
@@
Building on the work of Sesari et al. [4], who identified various scenarios in which fairness problems persist within SE contexts, we sought to explore these scenarios further. In our survey, respondents were asked to rate the extent to which they have experienced fairness across 10 SE contexts, adapted from the study by Sesari et al. [4]. Our analysis found that most respondents experienced unfairness in the context of demand (e.g., pressing deadline), authorship (e.g., code attribution), income (e.g., being underpaid), and allocation of work (e.g., assigned tasks unsuited to skills).
@@
III Background
III-A Developer Job Satisfaction
@@
In our study, we draw from Storey et al.’s understanding of job satisfaction [12], which is rooted in Wright and Cropanzano’s definition [15] who described job satisfaction as an internal state where an individual evaluates their job experiences positively or negatively.
@@
Beyond job satisfaction, developer motivation is another important factor frequently discussed in the literature. Beecham et al. reviewed SE studies and identified several factors influencing motivation, along with signs of both motivation and demotivation [16]. Following this, Sharp et al. proposed a framework that includes motivators, outcomes, and the context of software engineers’ work [17]. More recently, França et al. pointed out that factors such as career growth and autonomy are key to motivation, while also clarifying that motivation and job satisfaction are related [18, 19, 20, 21].
@@
Drawing inspiration from this research, Storey et al. incorporated relevant motivational factors into their study of developer satisfaction [12]. Storey et al. identified numerous social and technical factors, such as appreciation, rewards, work culture, and work-life balance, that influence job satisfaction in software development [12]. They also developed a theory linking higher job satisfaction to increased perceived productivity among software developers.
@@
In designing our survey, we took inspiration from their work, particularly the 44 factors they identified as influencing job satisfaction among software practitioners [12]. While their research focused broadly on these factors, our study seeks to explore fairness as a critical aspect, aiming to understand how fairness impacts job satisfaction in SE.
III-B Fairness in Software Engineering
@@
In this research, we define fairness in alignment with the broader justice literature. Fairness, as explained by Colquitt et al. [22] and Cuguero et al. [23], is a subjective assessment of whether organizational activities are appropriate or morally sound. Colquitt’s fairness theory divided fairness into several dimensions relevant to workplaces [5]. Distributive fairness addresses the fairness of outcomes, like rewards and recognition, and how well these align with principles such as equity or equality [24, 25, 26]. Procedural fairness concerns the fairness of the decision making processes, focusing on consistency, bias suppression, and opportunities for participation in decision-making [27, 28]. Later work introduced interpersonal fairness, which refers to respectful and dignified treatment during interactions with leads or managers, and informational fairness, which focuses on transparency and the clear communication of decisions with leads or managers [29].
@@
In the context of SE, fairness is increasingly seen as important in software development. German et al. [6] used Colquitt’s fairness theory [5] to develop a framework for making code reviews fairer. They highlighted the need for distributive fairness, which means reviews should consider the value of patches and decide whether to prioritize based on equity (i.e., importance of the contribution) or equality (i.e., treating all patches the same) [6]. This includes setting basic standards for patch reviews and making sure newcomers are treated fairly.
@@
For procedural fairness, the authors stressed the importance of having clear and consistent guidelines so that reviews are done without bias [6]. Procedural fairness also involves giving the author a chance to respond to feedback and creating an appeals process for rejected patches that involves new, unbiased reviewers. Interpersonal fairness is supported by codes of conduct, which aim to ensure respectful communication during reviews [6]. These codes should also protect authors’ privacy by preventing personal information from being used improperly. Informational fairness is maintained through the openness of code reviews in open-source projects [6]. Reviewers should give clear reasons when they reject patches and inform authors if a review is delayed.
@@
Sesari et al. further explored fairness concerns in SE through a review of developer posts on Stack Exchange sites, identifying key contexts where fairness issues arise [4]. These contexts include income, treatment, allocation of work, recruitment, working hours, demand, evaluation, authorship, policy, and customer/client relations. Their findings showed that most concerns focused on procedural fairness, such as work allocation process that is applied consistently, interview procedure and performance review that is based on accurate information, and software testing that is free from confirmation bias [4].
@@
Interestingly, while Sesari et al. found that many of the fairness issues are not exclusively tied to protected attributes such as gender, gender-related concerns remain a frequently discussed topic in fairness-related posts [4]. These studies suggest that fairness issues in SE extend beyond technical processes, affect the broader social dynamics within development teams both as a collective and from different backgrounds.
III-C How Fairness and Job Satisfaction are Related
@@
The relationship between fairness and job satisfaction has been extensively studied in organizational justice literature. Research has consistently shown strong links between procedural fairness and overall job satisfaction [30, 31]. Additionally, Masterson et al. found that procedural fairness was a stronger predictor of job satisfaction than interpersonal and informational fairness, though all were significant factors [32]. To explore the broader impact of fairness in organizational settings, Colquitt’s meta-analysis of organizational justice research examined the outcomes of 183 empirical studies across multiple fields, confirming the strong influence of perceived fairness on factors like organizational commitment, citizenship behaviors, and job satisfaction [7].
@@
Fairness has been shown to have a major impact on job satisfaction across various industries. For instance, fairness in management practices significantly influences nurse retention and job satisfaction in healthcare [8]. In the manufacturing sector, procedural fairness enhances employee commitment and decreases turnover intentions [9]. Similarly, in education, fairness in workload distribution and rewards strongly predicts teacher satisfaction [10].
@@
In the realm of SE, job satisfaction plays an essential role in improving employee retention, well-being, and productivity [11]. Many researchers and industry leaders assume that higher job satisfaction is closely tied to improved productivity among developers [12]. Storey et al. developed a theory linking increased job satisfaction to greater perceived productivity, identifying key factors like rewards, work culture, and work-life balance as crucial elements influencing satisfaction [12]. Developers in agile teams report higher job satisfaction than those in non-agile environments due to increased autonomy, decision-making power, and involvement in projects [33]. These factors relate to procedural fairness, as agile environments provide transparency and involvement in decision-making, making developers feel their contributions are fairly recognized.
@@
Personality traits like agreeableness and conscientiousness also influence job satisfaction [34]. Teams that allow developers to organize their work and minimize conflicts promote interpersonal fairness through respect and collaboration [34].Psychological safety and clear team norms are also found to be key predictors of job satisfaction, which allow developers to voice concerns without fear of judgment [35]. This fosters interpersonal fairness by creating a respectful environment, and informational fairness through well-communicated expectations, both of which contribute to job satisfaction. Building on this, our research explores whether perceptions of fairness are a key factor in developer job satisfaction, alongside other well-established factors.
IV Survey Design

### v2

# It Is Giving Major Satisfaction: Why Fairness Matters for Software Practitioners

@@
Abstract.
@@
Software practitioners often encounter workplace unfairness, such as unequal recognition and gender bias. While the link between fairness and job satisfaction has been established in other fields, its relevance to software professionals remains underexplored. This study examines how fairness perceptions relate to job satisfaction among software practitioners, focusing on both general trends and demographic-specific differences. We conducted an online survey of 108 software practitioners, followed by ordinal logistic regression to analyze the relationship between fairness perceptions and job satisfaction in software engineering contexts, with moderation analysis examining how this relationship varies across demographic groups. Our findings indicate that all four fairness dimensions (namely distributive, procedural, interpersonal, and informational fairness) significantly affect overall job satisfaction and satisfaction with job security. Among these, interpersonal fairness has the biggest impact. The relationship between fairness and job satisfaction is stronger for female, ethnically underrepresented, less experienced practitioners, and those with work limitations. Fairness in authorship emerged as an important factor for job satisfaction collectively, while fairness in policy implementation, high-demand situations, and working hours impacted specific demographic groups. This study highlights the role of fairness among software practitioners, offering strategies for organizations to promote fair practices and targeted approaches for certain demographic groups.
Fairness, job satisfaction, human and social aspects, software engineering
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Software and its engineering Collaboration in software developmentccs: Social and professional topics Socio-technical systems
1. Introduction
@@
In the field of Software Engineering (SE), practitioners frequently encounter fairness problems in various aspects of their work, such as problems in recognition of contributions (Casari et al., 2021), gender bias (Canedo et al., 2019), and ethnicity bias (Nadri et al., 2021). A study by Sesari et al.(Sesari et al., 2024) explored these fairness problems by identifying different contexts in which they arise, based on the well-established Colquitt’s organizational fairness framework (Colquitt, 2001). By leveraging this framework, the authors were able to pinpoint specific areas within SE where fairness issues are most prevalent. The top five areas identified were income, treatment, allocation of work, recruitment, and working hours (Sesari et al., 2024).
@@
Leveraging the same framework, German et al. (German et al., 2018) developed a framework explaining fairness in the modern code review process to identify strengths and potential areas for improvement. For example, distributive fairness, which includes principles like equity (Colquitt, 2001), suggests that decisions about how work is evaluated and rewarded should consider both the person who made the contribution and the quality of that contribution. German et al. (German et al., 2018) further studied fairness problems among developers involved in the OpenStack project by surveying them about the code review process. The survey results showed that many developers were concerned about fairness, especially regarding issues like equity, equal treatment, consistency, and preventing bias (German et al., 2018). However, while these problems had been identified, the impact of facing such fairness issues on software developers remains unclear.
@@
To better understand the broader implications of fairness in the workplace, Colquitt conducted a meta-analytic review of organizational justice literature, which included 120 separate meta-analyses of 183 empirical studies (Colquitt et al., 2001). This comprehensive review examined the effects of organizational fairness—encompassing distributive, procedural, interpersonal, and informational fairness—on several outcomes, including job satisfaction. Colquitt’s findings demonstrate a significant link between perceptions of fairness and job satisfaction across diverse contexts. For example, in the healthcare sector, fairness in managerial practices has been shown to significantly influence nurse retention and job satisfaction (Judge and Colquitt, 2004). In the manufacturing industry, procedural justice has been found to enhance employee commitment and reduce turnover intentions (Konovsky and Cropanzano, 1991). Similarly, in educational settings, distributive fairness related to workload and rewards is a key predictor of teacher satisfaction (Moorman, 1991). Given these findings across various fields, an important question arises: does the observed link between fairness and job satisfaction apply in the context of SE?
@@
In the SE literature, enhanced job satisfaction among developers is known to improve employee attraction, retention, and well-being (Fagerholm et al., 2014). Furthermore, many researchers and companies assume that more satisfied developers tend to be more productive (Storey et al., 2021). Storey et al. (Storey et al., 2021) developed a theory that links higher job satisfaction among software developers to increased perceived productivity. In their study, they also identified various social and technical factors that significantly impact job satisfaction, such as appreciation and rewards, work culture, and work life balance.
@@
Building on this work, we are interested in exploring perception of fairness as another potential factor that could shape job satisfaction. Similar to Storey et al.’s (Storey et al., 2021) approach in identifying key factors that influence job satisfaction, our study seeks to investigate whether and how perceptions of fairness relate to job satisfaction among software practitioners, both as a collective, and from different backgrounds.
2. Research Questions
@@
To address whether and how perceptions of fairness relate to job satisfaction, we conducted an online survey, receiving a total of 108 responses. We used the 20-item fairness scale developed by Colquitt (Colquitt, 2001) to measure fairness perceptions across four dimensions: distributive, procedural, interpersonal, and informational fairness. Job satisfaction was measured using 15 satisfaction factors inspired by Storey et al. (Storey et al., 2021).
@@
This observation aligns with existing literature that highlights demographic disparities in the SE industry, which is predominantly young, White, and male. Studies have shown that minority groups often face unique challenges. For example, Trinkenreich et al. identified that women in the software industry encounter issues like work-life balance, the glass ceiling, and sexism (Trinkenreich et al., 2022a). Canedo et al. found that women are underrepresented in software development and open source projects, often due to male-dominated workplace conditions (Canedo et al., 2019). Furthermore, Nadri et al. reported that less than 10% of contributions to open source projects come from visibly Non-White developers (Nadri et al., 2021). Similarly, Baltes et al. discuss age-related biases, noting that older developers are often perceived as less adaptable, which can limit their career progression (Baltes et al., 2020).
@@
Building on the work of Sesari et al. (Sesari et al., 2024), who identified various scenarios in which fairness problems persist within SE contexts, we sought to explore these scenarios further. In our survey, respondents were asked to rate the extent to which they have experienced fairness across 10 SE contexts, adapted from the study by Sesari et al. (Sesari et al., 2024). Our analysis found that most respondents experienced unfairness in the context of demand (e.g., pressing deadline), authorship (e.g., code attribution), income (e.g., being underpaid), and allocation of work (e.g., assigned tasks unsuited to skills).
@@
3. Background
3.1. Developer Job Satisfaction
@@
In our study, we draw from Storey et al.’s understanding of job satisfaction (Storey et al., 2021), which is rooted in Wright and Cropanzano’s definition (Wright and Cropanzano, 2000) who described job satisfaction as an internal state where an individual evaluates their job experiences positively or negatively.
@@
Beyond job satisfaction, developer motivation is another important factor frequently discussed in the literature. Beecham et al. reviewed SE studies and identified several factors influencing motivation, along with signs of both motivation and demotivation (Beecham et al., 2008). Following this, Sharp et al. proposed a framework that includes motivators, outcomes, and the context of software engineers’ work (Sharp et al., 2009). More recently, França et al. pointed out that factors such as career growth and autonomy are key to motivation, while also clarifying that motivation and job satisfaction are related (França et al., 2012; Franca et al., 2013; França et al., 2014, 2018).
@@
Drawing inspiration from this research, Storey et al. incorporated relevant motivational factors into their study of developer satisfaction (Storey et al., 2021). Storey et al. identified numerous social and technical factors, such as appreciation, rewards, work culture, and work-life balance, that influence job satisfaction in software development (Storey et al., 2021). They also developed a theory linking higher job satisfaction to increased perceived productivity among software developers.
@@
In designing our survey, we took inspiration from their work, particularly the 44 factors they identified as influencing job satisfaction among software practitioners (Storey et al., 2021). While their research focused broadly on these factors, our study seeks to explore fairness as a critical aspect, aiming to understand how fairness impacts job satisfaction in SE.
3.2. Fairness in Software Engineering
@@
In this research, we define fairness in alignment with the broader justice literature. Fairness, as explained by Colquitt et al. (Colquitt and Rodell, 2015) and Cuguero et al. (Cugueró-Escofet and Rosanas, 2013), is a subjective assessment of whether organizational activities are appropriate or morally sound. Colquitt’s fairness theory divided fairness into several dimensions relevant to workplaces (Colquitt, 2001). Distributive fairness addresses the fairness of outcomes, like rewards and recognition, and how well these align with principles such as equity or equality (Adams, 1965; Deutsch, 1975; Leventhal, 1976). Procedural fairness concerns the fairness of the decision making processes, focusing on consistency, bias suppression, and opportunities for participation in decision-making (Leventhal, 1980; Thibaut and Walker, 1975). Later work introduced interpersonal fairness, which refers to respectful and dignified treatment during interactions with leads or managers, and informational fairness, which focuses on transparency and the clear communication of decisions with leads or managers (Greenberg, 1993).
@@
In the context of SE, fairness is increasingly seen as important in software development. German et al. (German et al., 2018) used Colquitt’s fairness theory (Colquitt, 2001) to develop a framework for making code reviews fairer. They highlighted the need for distributive fairness, which means reviews should consider the value of patches and decide whether to prioritize based on equity (i.e., importance of the contribution) or equality (i.e., treating all patches the same) (German et al., 2018). This includes setting basic standards for patch reviews and making sure newcomers are treated fairly.
@@
For procedural fairness, the authors stressed the importance of having clear and consistent guidelines so that reviews are done without bias (German et al., 2018). Procedural fairness also involves giving the author a chance to respond to feedback and creating an appeals process for rejected patches that involves new, unbiased reviewers. Interpersonal fairness is supported by codes of conduct, which aim to ensure respectful communication during reviews (German et al., 2018). These codes should also protect authors’ privacy by preventing personal information from being used improperly. Informational fairness is maintained through the openness of code reviews in open-source projects (German et al., 2018). Reviewers should give clear reasons when they reject patches and inform authors if a review is delayed.
@@
Sesari et al. further explored fairness concerns in SE through a review of developer posts on Stack Exchange sites, identifying key contexts where fairness issues arise (Sesari et al., 2024). These contexts include income, treatment, allocation of work, recruitment, working hours, demand, evaluation, authorship, policy, and customer/client relations. Their findings showed that most concerns focused on procedural fairness, such as work allocation process that is applied consistently, interview procedure and performance review that is based on accurate information, and software testing that is free from confirmation bias (Sesari et al., 2024).
@@
Interestingly, while Sesari et al. found that many of the fairness issues are not exclusively tied to protected attributes such as gender, gender-related concerns remain a frequently discussed topic in fairness-related posts (Sesari et al., 2024). These studies suggest that fairness issues in SE extend beyond technical processes, affect the broader social dynamics within development teams both as a collective and from different backgrounds.
3.3. How Fairness and Job Satisfaction are Related
@@
The relationship between fairness and job satisfaction has been extensively studied in organizational justice literature. Research has consistently shown strong links between procedural fairness and overall job satisfaction (Mossholder et al., 1998; Wesolowski and Mossholder, 1997). Additionally, Masterson et al. found that procedural fairness was a stronger predictor of job satisfaction than interpersonal and informational fairness, though all were significant factors (Masterson et al., 2000). To explore the broader impact of fairness in organizational settings, Colquitt’s meta-analysis of organizational justice research examined the outcomes of 183 empirical studies across multiple fields, confirming the strong influence of perceived fairness on factors like organizational commitment, citizenship behaviors, and job satisfaction (Colquitt et al., 2001).
@@
Fairness has been shown to have a major impact on job satisfaction across various industries. For instance, fairness in management practices significantly influences nurse retention and job satisfaction in healthcare (Judge and Colquitt, 2004). In the manufacturing sector, procedural fairness enhances employee commitment and decreases turnover intentions (Konovsky and Cropanzano, 1991). Similarly, in education, fairness in workload distribution and rewards strongly predicts teacher satisfaction (Moorman, 1991).
@@
In the realm of SE, job satisfaction plays an essential role in improving employee retention, well-being, and productivity (Fagerholm et al., 2014). Many researchers and industry leaders assume that higher job satisfaction is closely tied to improved productivity among developers (Storey et al., 2021). Storey et al. developed a theory linking increased job satisfaction to greater perceived productivity, identifying key factors like rewards, work culture, and work-life balance as crucial elements influencing satisfaction (Storey et al., 2021). Developers in agile teams report higher job satisfaction than those in non-agile environments due to increased autonomy, decision-making power, and involvement in projects (Melnik and Maurer, 2006). These factors relate to procedural fairness, as agile environments provide transparency and involvement in decision-making, making developers feel their contributions are fairly recognized.
@@
Personality traits like agreeableness and conscientiousness also influence job satisfaction (Acuña et al., 2009). Teams that allow developers to organize their work and minimize conflicts promote interpersonal fairness through respect and collaboration (Acuña et al., 2009). Psychological safety and clear team norms are also found to be key predictors of job satisfaction, which allow developers to voice concerns without fear of judgment (Lenberg and Feldt, 2018). This fosters interpersonal fairness by creating a respectful environment, and informational fairness through well-communicated expectations, both of which contribute to job satisfaction. Building on this, our research explores whether perceptions of fairness are a key factor in developer job satisfaction, alongside other well-established factors.
4. Survey Design